OpenStudy (anonymous):

Peter sold two flats for $999999 each. He lost 10% on one and gained 10% on the other. After the two transactions, Peter: A. gained $10101 B. gained$20202 C. lost $10101 D. lost $20202

OpenStudy (phi):

sold two flats for $999999 each that means get received 999,999 on the one he lost 10% 0.9*x = 999,999 we can can find x. x= 999,999/0.9 he lost 10% of x 0.1*x = 0.1/0.9 * 999,999 = 1/9 * 999,999 = 111,111 lost

OpenStudy (anonymous):

\[ 999999\times (1.1)-999999\times .9\]

OpenStudy (anonymous):

but there're only two values for us to choose: 10101 and 20202

OpenStudy (phi):

he received 999,999 on the one he made 10% 1.1x= 999,999 x= 999,999/1.1 he made 10% of this 0.1 x = 999,999/11 = 90,909 gained

OpenStudy (anonymous):

ohh

OpenStudy (phi):

111,111 lost + 90,909 gained or lost 20,202

OpenStudy (anonymous):

thz